calculate normality and titre of koh -solution (volume 5.45 ml) if it is necessary for the titration 6.7 ml h2so4-solution with cn=0.1025 mol/l


1
Expert's answer
2021-11-24T12:21:16-0500

Solution:

CN(KOH) × V(KOH) = CN(H2SO4) × V(H2SO4)

Therefore,

CN(KOH) = CN(H2SO4) × V(H2SO4) / V(KOH)

CN(KOH) = (0.1025 mol L-1 × 6.7 mL) / (5.45 mL) = 0.1260 mol L-1

Normality of KOH solution is 0.1260 mol L-1


CN(KOH) = CM(KOH)

Molarity of KOH solution is 0.1260 mol L-1


The molar mass of KOH is 56.1 g mol-1

Therefore the titre is:

(0.1260 mol L-1) × (56.1 g mol-1) = 7.0686 g L-1 = 7.0686 mg mL-1

Titre of KOH solution is 7.0686 mg mL-1


Answer:

Normality of KOH solution is 0.1260 mol L-1

Titre of KOH solution is 7.0686 mg mL-1